OpenStudy (anonymous):

Determine whether the sequence could be arithmetic. If so, find the common difference and the next term.

OpenStudy (anonymous):

Take the difference between consecutive terms and see if it is the same throughout the sequence.

OpenStudy (anonymous):

\[\frac{ 11 }{ 2 },\frac{ 11 }{ 3 },\frac{ 11 }{ 4 },\frac{ 11 }{ 5 },\frac{ 11 }{ 6 }\]

OpenStudy (loser66):

@Mertsj

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

I'll use alternate terms: 11/2-11/4=11/4. If this were an arithmetic sequence, the 5th term would be 11/4-11/4=0. It is not, therefore this is not an arithmetic sequence.

OpenStudy (anonymous):

@Luis_Rivera

OpenStudy (anonymous):

@Luis_Rivera can you help?

OpenStudy (anonymous):

It isn't. It's a geometric relationship, but not a geometric sequence. If it were a geometric sequence, instead of the 1/(n+1) you would have a simple exponential of the ratio r.

OpenStudy (anonymous):

so its not arithmetic?

OpenStudy (anonymous):



OpenStudy (anonymous):

uhh, yes it is geometic

OpenStudy (anonymous):

It is neither arithmetic nor geometric...

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

Arithmetic sequences can be written in the format:\[a_n=a_0+n \times c\]Geometric sequences can be written in the format:\[a_n=a_0 \times r^n\]This sequence is neither.

OpenStudy (anonymous):

For the love of... this sequence has a formula:\[a_n=a_0 \div (n+1)\] which is neither of the formulas above, so it's not arithmetic or geometric.

OpenStudy (anonymous):

\[uhh, \ \ a _{0} \div{n+1}= a _{0}*\frac{ 1 }{ n+1 }\ \ \ so \ its \ geometric\]

OpenStudy (anonymous):

\[a_{n}=a _{1}+(n+1)d\]

OpenStudy (anonymous):

@Numb3r1 @Luis_Rivera

OpenStudy (anonymous):

thats the real formula

OpenStudy (anonymous):

That's not how geometric sequences work, @Luis_Rivera. @tomtom777, your formula is not the neatest formula for an arithmetic sequence, but it works. However, this is NOT AN ARITHMETIC OR GEOMETRIC SEQUENCE. Once more: Arithmetic sequences can be written in the format: an=a0+n×c Geometric sequences can be written in the format: an=a0×r^n This sequence is neither: a0÷(n+1)=a0∗(1÷(n+1)) is NOT ONE OF THE ABOVE FORMULAS. Ignore both of them, they are simply wrong.